OpenStudy (anonymous):

How do I solve log3x+log8=3? round to the nearest hundreath if necessary

OpenStudy (agent0smith):

use this loga + logb = log(ab) on this: log3x+log8

OpenStudy (anonymous):

log 3x*8=3 3x*8=antilog(3) 3x*8=1000 x=1000/24